### v1

# Universal scaling limits for spin networks via martingale methods

Abstract
We use simple martingale methods to construct a large deviation theory of spin systems with pairwise interactions. As an application, we show that the fully connected case obeys a universal scaling limit that is just a product of magnetisation eigenstates.
Keywords: networks, generative models, graphs, spin systems
Sapienza Università di Roma, Piazza Aldo Moro 1, 00185 Roma, Italy
In [1, 2, 3] we elaborated a large–deviations theory for disordered systems that allows to study any spin systems by means of simple combinatorial techniques, including the celebrated Sherrington–Kirckpatrick (SK) model, a mean field paradigm for complex systems [4]. Here, we used the same methods to show that any fully connected models with any pairwise interaction, including the random models (i.e., models with random parameters like the SK model), obeys a universal scaling limit. We actually construct the scaling limit of the free energy density in terms of magnetisations eigenstates, i.e., spin states with fixed average magnetisation, in the Eq. (85).
The main points of our analysis where to introduce a sequence of sub–systems of increasing size by partitioning the vertex set into a number of disjoint sub–sets (that in the references [1, 2, 3] we called “layers”) and then by re–joining them up to a certain level of that sequence: this is shown in the Section 1. Then, we reinterpret this as a series of Markov blankets, [5, 6], and identify their Hamiltonian sequence in Section 2. We find that each blanket can be further destructured in two general contributions: the first term, “core” contribution, is the energy from the interactions inside the blanket itself, while the second, the “interface”, is due to the interactions between the spins of the blanket and the rest of the network. Formally, it is the scalar product between the spins of the blanket and local fields that depends on the previous blankets. As we shall see this is a general theory, and up to this point it could be used to deal with any spin system. Then, in Section 3, we show that in fully connected models if the partition is fine enough the core contribution is small, and the thermodynamics is dominated by the interfaces. Finally, in the Section 3.1, we show that the Hamiltonian with the cores removed (IO model, [3]) obeys a universal scaling limit that can be expressed in terms of the magnetisation eigenstates like in the SK model [3, 7].
Apart from providing characterisation for the scaling limit of any fully connected spin network in terms of magnetisation eigenstates, an important goal of this paper is to provide a simple presentation of the method introduced in [1, 2, 3] that is readable by people working in the fields of Bayesian Inference, Machine Learning [5], and Active Inference [6, 8, 9, 10]. Given this purpose, here we avoid the details of the underlying Kernel Theory (KT) [3, 10, 11, 19, 20], within which these results where obtained. The interested reader may look at the Sections 2 and 3 of [3], Section 4 of [10] the SI section of [11] and the preprints [19, 20] for a comprehensive presentation.
Notation and fundamentals
Let us introduce the notation. The set of vertices on which the spins are placed will be denoted by $V$, with the vertices tracked by the integer label $i$
$V:=\{1\leq i\leq N\},$ (1)
the total number of vertices will be denoted by $N$, which from here on is assumed to be a thermodynamically large integer. The binary support of a single spin is
$\Omega:=\{-1,1\}$ (2)
The generic spin state (magnetisation state) is as follows
$\sigma_{V}:=\{\sigma_{i}\in\Omega:\,i\in V\}\in\Omega^{V}$ (3)
and represents one of the possible configurations of the system. Since the support of the single spin is limited to two states, the cardinality of the support of the magnetisations states will be equal to $|\Omega|=2$ raised to the number of vertices $N$,
$\left|\Omega^{V}\right|=2^{N}$ (4)
in general, when we apply the absolute value to a set we mean its cardinality. In what follows we will consider fixed the order $1\leq i\leq N$ in which the spins are labelled, this is reflected in the nonstandard notation of the support
$\Omega^{\{1\}}\,\Omega^{\{2\}}\ ...\ \ \Omega^{\{N\}}=\prod_{i\in V}\ \Omega^{\{i\}}=:\Omega^{V}$ (5)
where, instead of simply raising to the power $N$ the generic binary support $\Omega$, here we also want to keep track of the order in which the individual spin
$\Omega^{\{i\}}\ni\sigma_{i}$ (6)
are ordered to obtain the full magnetisation state $\sigma_{V}$. It will be convenient to introduce a special notation for the uniform measure: the uniform average will be indicated with the angled brackets with the support in subscript
$\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{\Omega^{V}}:=\frac{1}{|\,\Omega^{V}|}\sum_{\sigma_{V}\in\Omega^{V}}\mathcal{O}\left(\sigma_{V}\right)$ (7)
We also introduce the eigenstates of magnetisation, which we might regard as the fundamental building blocks of our description. We call ‘total magnetisation’ the operator that, if applied to a spin state, returns its total magnetisation. The ‘eigenstates of magnetization’ with eigenvalue $M$ are therefore those spin states with magnetisation exactly equal to $M$. Then, let us introduce the magnetisation function
$M(\sigma_{V}):=\sum_{i\in V}\sigma_{i},\ \ \ m(\sigma_{V}):=\frac{M(\sigma_{V})}{N},$ (8)
that tracks the magnetisation of the spin vector. The value of the magnetisation varies from the maximum $N$ to the minimum $-N$, corresponding to all spin up and all spin down, respectively. Labelling with this parameter we can further partition the support of $\sigma_{V}$ into disjoint subsets of given magnetisation:
$\Omega^{V}\left(M\right):=\{\sigma_{V}\in\Omega^{V}:\,M(\sigma_{V})=M\}$ (9)
the set collects all states of size $N$ and total magnetisation $M$. The complete support is obtained by joining together the supports with given magnetisation
$\Omega^{V}=\bigcup_{M}\ \Omega^{V}\left(M\right)$ (10)
As for the case of the uniform average, we will use the angle brackets (with the support in subscript) to indicate the average respect to the magnetisation eigenstates,
$\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{\,\Omega^{V}\left(M\right)}:=\frac{1}{|\,\Omega^{V}\left(M\right)|}\sum_{\sigma_{V}\in\Omega^{V}\left(M\right)}\mathcal{O}\left(\sigma_{V}\right)$ (11)
Magnetisation eigenstates exist also in the thermodynamic limit. To see this, let
$M=\left\lfloor mN\right\rfloor,\ \ \ m\in\left[-1,1\,\right].$ (12)
From the theory of large deviations it is possible to deduce various properties of eigenstates in the continuous limit as the magnetisation parameter $m$ varies. To lighten the notation we introduce the following abbreviations.
$\Omega^{V}\left(m\right):=\Omega^{V}\left(\left\lfloor mN\right\rfloor\right)$ (13)
The average over the eigenstates with parameter $\left\lfloor mN\right\rfloor$ will be denoted simply by
$\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{m}:=\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{\,\Omega^{V}\left(\left\lfloor mN\right\rfloor\right)}$ (14)
Since the eigenstates of the magnetisation are an ergodic set, the average on the eigenstates can be computed numerically by generating samples with a given magnetisation, evaluating the function and ultimately averaging. Given the binary nature of the magnetisation values for the single spin, it is also possible to obtain a detailed description in the thermodynamic limit at the sample–path level (sample–path Large Deviations) by applying the Varadhan lemma and the Mogulskii theorem. For the teaching purpose of this paper we will not go into these more advanced methods, but the interested reader may consult [12, 13, 14] and references therein for more details.
1 The general model
The general model we are going to study is a simple binary field theory, truncated at second-order. For simplicity let us consider an isolated system (with no external field), given an arbitrary interaction matrix
$H=\{H_{ij}\in\Gamma:\,ij\in V^{2}\}\in\Gamma^{\,V^{2}}$ (15)
where the elementary support $\Gamma$ can be a finite alphabet or interval or directly the real axis. The Hamiltonian of the model is in general
$H\left(\sigma_{V}\right):=\sum_{i\in V}\sum_{j\in V}H_{ij}\sigma_{i}\sigma_{j}$ (16)
this can be rewritten in a way that is convenient for us by introducing the local fields, or cavity fields, depending on the context. By grouping the local fields into a vector
$h_{V}\left(\sigma_{V}\right):=\{h_{i}\left(\sigma_{V}\right)\in\mathbb{R}:\,i\in V\},\ \ \ h_{i}\left(\sigma_{V}\right):=\sum_{j\in V}H_{ij}\sigma_{j}$ (17)
the Hamiltonian is rewritten as product between the state and the local fields
$H\left(\sigma_{V}\right)=\sigma_{V}\cdot h_{V}\left(\sigma_{V}\right).$ (18)
The partition function and the associated Gibbs measure (softmax) are respectively
$Z:=\sum_{\sigma_{V}\in\Omega^{V}}\exp\left[-H\left(\sigma_{V}\right)\right],\ \ \ \mu\left(\sigma_{V}\right):=\frac{\exp\left[-H\left(\sigma_{V}\right)\right]}{Z}$ (19)
we will use the symbol $\sim$ to indicate that a variable is distributed according to a certain measure, and the symbol $\mathcal{P}$ to indicate the probability space generated by the support within parentheses, for example we will write
$\sigma_{V}\sim\mu\in\mathcal{P}(\Omega^{V})$ (20)
to indicate that the magnetisation state is distributed according to the Gibbs measure $\mu$, which belongs to the set of possible probability measures defined on the support $\Omega^{V}$. The average respect to the generic probability measure will be indicated with angled parenthesis with the measure indicated in subscript, for example:
$\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{\mu}:=\sum_{\sigma_{V}\in\Omega^{V}}\mu\left(\sigma_{V}\right)\mathcal{O}\left(\sigma_{V}\right).$ (21)
Apart from the case of uniform measure and magnetisation eigenstates, this will be our standard notation for averaging over a given measurement.
1.1 Thermodynamic limit
In canonical statistical theory, the thermodynamics of the model is entirely described by the Helmholtz free energy function
$f:=-\frac{1}{N}\log Z$ (22)
The very existence of the thermodynamic limit can in fact be reduced to the existence of the limit of $f$ when $N\rightarrow\infty$, which depends critically on how $H_{ij}$ scales in the spin number. The most general form could be site–dependent, but, for simplicity, we will consider a smaller class in which the normalisation is uniform:
$H_{ij}=\frac{\beta J_{ij}}{g\left(V\right)}$ (23)
where $J_{ij}$ are the coupling parameters and $g\left(V\right)$ is a normalisation that ensures the existence of $f$ in the thermodynamic limit, ie. such that the components of $h_{V}$ converges to finite numbers (at least in distribution). For a fully connected model this is usually diverging in the spin number, and not dependent from the specific labels of the spins. For example in the fully connected Curie-Weiss model we find that the correct normalisation must scale with $N$, while for the version with random weights, that is the SK model, the correct normalisation is instead $\sqrt{N}$. Notice that to lighten the notation we reabsorbed also the temperature in the parameters $H_{ij}$.
1.2 Free energy principle
Applying the Jensen inequality we can easily deduce the canonical free energy principle [6, 8, 10]. Let define the free energy functional
$\mathscr{F}\left(\xi\right):=\langle H\left(\sigma_{V}\right)\rangle_{\xi}+\langle\,\log\xi\left(\sigma_{V}\right)\rangle_{\xi}$ (24)
and consider the following chain of inequalities [3, 8]
$\sum_{\sigma_{V}\in\Omega^{V}}\exp\left[-H\left(\sigma_{V}\right)\right]=\sum_{\sigma_{V}\in\Omega^{V}}\xi\left(\sigma_{V}\right)\exp\left[-H\left(\sigma_{V}\right)-\log\xi\left(\sigma_{V}\right)\right]=\\ =\langle\exp\left[-H\left(\sigma_{V}\right)-\log\xi\left(\sigma_{V}\right)\right]\rangle_{\xi}\geq\exp\left[-\mathscr{F}\left(\xi\right)\right]$ (25)
where in the last step we applied the Jensen inequality to the exponential to move the average inside. Then for any $\xi$ holds
$Nf\leq\mathscr{F}\left(\xi\right),\ \ \ \forall\xi\in\mathscr{P}(\Omega^{V})$ (26)
It can be shown that the minimum is the Helmholtz free energy,
$Nf=\inf_{\xi\in\mathscr{P}(\Omega^{V})}\mathscr{F}\left(\xi\right)$ (27)
that is attained by the Gibbs measure $\mu$. This was the canonical free energy principle. We can also deal with random Hamiltonians. Let now suppose that the coupling matrix is distributed according to some probability measure $\eta$
$H\sim\eta\in\mathcal{P}(\Gamma^{\,V^{2}})$ (28)
and that we are interested in computing the quenched free energy,
$\langle f\rangle_{\eta}:=\sum_{H\in\Gamma^{\,V^{2}}}\eta f$ (29)
Applying Jensen inequality again (to the logarithm this time) we find the corresponding lower variational bound. From definitions we immediately obtain
$-N\langle f\rangle_{\eta}=\sum_{H\in\Gamma^{\,V^{2}}}\eta\log Z=\sum_{H\in\Gamma^{\,V^{2}}}\log Z^{\,\eta}=\\ =\sum_{H\in\Gamma^{\,V^{2}}}\zeta\log Z^{\,\eta/\zeta}=\langle\log Z^{\,\eta/\zeta}\rangle_{\zeta},\ \ \ \forall\zeta\in\mathscr{P}(\Gamma^{\,V^{2}})$ (30)
Let introduce the quenched free energy functional
$\mathscr{G}\left(\zeta\right):=-\log\,\langle\,Z^{\,\eta/\zeta}\rangle_{\zeta}=-\log\,\langle\exp\,\left(-N\eta f/\zeta\right)\rangle_{\zeta}$ (31)
By Jensen inequality holds the following
$N\langle f\rangle_{\eta}\geq\mathscr{G}\left(\zeta\right),\ \ \ \forall\zeta\in\mathscr{P}(\Gamma^{\,V^{2}})$ (32)
from which we deduce a variational principle for the quenched free energy
$N\langle f\rangle_{\eta}=\sup_{\zeta\in\mathscr{P}(\Gamma^{\,V^{2}})}\mathscr{G}\left(\xi\right)$ (33)
that is attained by the following measure
$\rho:=\frac{\eta f}{\langle f\rangle_{\eta}}$ (34)
this principle that takes the supremum of the functional instead of the infimum is reminiscent of the Parisi variational principle for the SK model or the Guerra interpolation. Notice that both principles where obtained by inserting the identity operator in the form of either $\log\,\exp$ or $\exp\log$ and then bringing the average one step inside,
$\langle f\rangle=\langle\exp\,\log\left(f\right)\rangle\geq\exp\,\langle\log\left(f\right)\rangle,\ \ \ \langle f\rangle=\langle\log\,\exp\left(f\right)\rangle\leq\log\,\langle\exp\left(f\right)\rangle$ (35)
and should be ultimately equivalent. Notice also that combining with the free energy principle we can introduce one last functional
$-\mathscr{A}\left(\xi,\zeta\right):=\log\,\langle\exp\left[-\eta\mathscr{F}\left(\xi\right)/\zeta\right]\rangle_{\zeta}=\\ =\log\,\langle\exp\,[-\eta\langle H\rangle_{\xi}/\zeta-\eta\langle\log\xi\rangle_{\xi}/\zeta]\rangle_{\zeta}$ (36)
that allows to write the free energy in terms of a min-max principle
$\mathscr{G}\left(\zeta\right)=\inf_{\xi\in\mathscr{P}(\Omega^{V})}\mathscr{A}\left(\xi,\zeta\right)$ (37)
This is reminiscent of some methods introduced by Guerra et al. and may be related to the min-max principle identified by these authors for the bipartite SK model [15].
2 Blanket (or Layer) representation
Now that we have introduced the fundamental quantities we can summarise the method proposed in [3]. The theory builds on several Graph–Theoretic methods for statistical physics and probability theory introduced by Lovaz, Borgs, Chayes, Coja-Oghlan, and others, [16, 17, 18, 19, 20]. The idea can be understood also in terms of partitioning the spins into a progression of Markov blankets [5, 6] (that in [1, 2, 3] we called “layers”). For now we will not specify the detailed structure of the blankets, which in general depends on the model, we will just discuss the most general features. Then, let introduce a partition of $V$ into $L$ disjoint subsets, hereafter recalled by the symbol $\mathscr{V}$,
$V=\left\{V_{\ell}\in\mathscr{V}:\,\ell\in\Lambda\right\},\ \ \ \Lambda:=\left\{1\leq\ell\leq L\right\}.$ (38)
As we shall see in short, the choice of the partition will be our general order parameter. To facilitate comparison with other methods it will be convenient to introduce a nested sequence of sets constructed by progressively bringing the parts together following the order established by the index $\ell$, then we introduce
$Q_{\ell}:=\bigcup_{\ell^{*}\leq\ell}V_{\ell^{*}},\ \ \ Q_{\ell-1}\subseteq Q_{\ell},\ \ \ Q_{L}=V,$ (39)
This sequence is completely determined by the partition $\mathscr{V}$,
$V_{\ell}:=Q_{\ell}\setminus Q_{\ell-1}.$ (40)
As we shall see, for densely connected models like the SK model only the relative size of the parts will matter, then we introduce the functional order parameter
$q:=\left\{q_{\ell}\in\left[0,1\right]:\,q_{0}=0,\,q_{L}=1,\,q_{\ell}\leq q_{\ell+1},\,\ell\in\Lambda\right\}$ (41)
that is a non–decreasing step function of $L$ steps bounded between zero and one. The definition implies that $q$ is the quantile function [21] of some atomic probability measure. In fact, the individual atoms, or steps
$p_{\ell}:=q_{\ell}-q_{\ell-1}\in\left[0,1\right]$ (42)
can be actually interpreted as probabilities. Then, the order parameter is also uniquely associated to an atomic probability distribution
$p:=\left\{p_{\ell}\in\left[0,1\right]:\,\ell\in\Lambda\right\}\in\mathscr{P}(\Lambda).$ (43)
and we could equivalently consider the distribution $p$ as order parameter. The relation between the size of the parts and the number of spins is
$|Q_{\ell}|=Nq_{\ell},\ \ \ |V_{\ell}|=N\left(q_{\ell}-q_{\ell-1}\right)=Np_{\ell}.$ (44)
Notice that the partition of $V$ induces a corresponding partition $\mathscr{W}\left(\mathscr{V}\right)$ in the edges set $V^{2}$ that allows to chart the energy contributions in the Hamiltonian. Let introduce
$W_{\ell}:=Q_{\ell}^{2}\setminus Q_{\ell-1}^{2},$ (45)
that contains all the edges with both ends in $Q_{\ell}$ minus those with both ends in $Q_{\ell-1}$. The cardinalities of these sets are linked to the order parameter by
$|W_{\ell}|=|Q_{\ell}^{2}|-|Q_{\ell-1}^{2}|=N^{2}(\,q_{\ell}^{2}-q_{\ell-1}^{2})=N^{2}(p_{\ell}^{2}+2\,q_{\ell}p_{\ell})$ (46)
These coefficients are known to Spin Glass theory since they appear in the Parisi functional for the SK model. Notice that $p$ is the discrete canonical conjugate of $q$, and that $p_{\ell}^{2}$ is an infinitesimal quantity that can be ignored in the large $L$ limit. Then, in the large $L$ limit the coefficient converges to the derivative of $q_{\ell}^{2}$ respect to $\ell$.
Figure 1: A comparison between the notion of a Markov blanket as is usually intended (e.g., by Pearl [5] and Friston [6]) versus the ‘layer’ representation of [1, 2, 3]. As one can see from the figures, the second case is just the interaction picture of the first: the blanket is made of vertices while the layer is made of edges. For this paper we will not make the distinction and refer to both as Markov blankets. This is done to facilitate the understanding with other fields that already use this concept.
2.1 Blankets and conditional expectations
We can now apply the partition to the energy function and see if this help in some way. Doing this, in [3] we identify the Hamiltonian
$H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right):=\sum_{\left(i,j\right)\in W_{\ell}}H_{ij}\sigma_{i}\sigma_{j}$ (47)
as the energy function that describes the particular blanket $\ell$. The full Hamiltonian is recovered by simply summing the blankets
$H\left(\sigma_{V}\right)=\sum_{\ell\leq L}H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)$ (48)
For each blanket we can associate the partition function
$Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right):=\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\exp\left[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\right]$ (49)
and the associated Gibbs measure of the blanket
$\xi_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right):=\frac{\exp\left[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\right]}{Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)}$ (50)
The full probability distribution and the free energy functional are obtained by multiplying the Gibbs measures of the individual blankets,
$\mu\left(\sigma_{V}\right):=\prod_{\ell\leq L}\xi_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right),\ \ \ Z\left(\sigma_{V}\right):=\prod_{\ell\leq L}Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)$ (51)
In particular, the second functional in the previous equation is related to the partition function by the following fundamental formula:
$Z=\langle Z\left(\sigma_{V}\right)\rangle_{\mu},\ \ \ \forall p\in\mathscr{P}(\Lambda)$ (52)
that express the flatness of the functional in Eq. (51) respect to variations of the order parameter $p$. This result can be shown easily by the following chain of identities:
$Z=\sum_{\sigma_{V_{1}}\in\Omega^{V_{1}}}\,...\,\sum_{\sigma_{V_{L}}\in\Omega^{V_{L}}}\ \prod_{\ell\leq L}\ \exp\,[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)]=\\ =\sum_{\sigma_{V_{1}}\in\Omega^{V_{1}}}\exp\left[-H_{1}\left(\sigma_{V_{1}}\right)\right]\ \ \ ...\sum_{\sigma_{V_{L}}\in\Omega^{V_{L}}}\exp\left[-H_{L}\left(\sigma_{V_{L}}|\sigma_{Q_{L-1}}\right)\right]=\\ =\sum_{\sigma_{V_{1}}\in\Omega^{V_{1}}}Z_{1}\,\xi_{1}\left(\sigma_{V_{1}}\right)\ \ \ ...\sum_{\sigma_{V_{L}}\in\Omega^{V_{L}}}Z_{L}\left(\sigma_{Q_{L-1}}\right)\,\xi_{L}\left(\sigma_{V_{L}}|\sigma_{Q_{L-1}}\right)=\\ =\sum_{\sigma_{V_{1}}\in\Omega^{V_{1}}}\xi_{1}\left(\sigma_{V_{1}}\right)\ \ \ ...\sum_{\sigma_{V_{L-1}}\in\Omega^{V_{L-1}}}\xi_{L-1}\left(\sigma_{V_{L-1}}|\sigma_{Q_{L-2}}\right)\prod_{\ell\leq L}Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=\\ =\sum_{\sigma_{V_{1}}\in\Omega^{V_{1}}}\ \ \ ...\sum_{\sigma_{V_{L-1}}\in\Omega^{V_{L-1}}}\,\prod_{\ell<L}\xi_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\prod_{\ell\leq L}Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=\\ =\sum_{\sigma_{Q_{L-1}}\in\Omega^{Q_{L-1}}}\,\prod_{\ell<L}\xi_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\prod_{\ell\leq L}Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=\\ =\langle\prod_{\ell\leq L}Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)\rangle_{\mu}=\langle Z\left(\sigma_{V}\right)\rangle_{\mu}.$ (53)
In general, the average of any observable $\mathcal{O}$ according to $\mu$ is found trough a cascade application of conditional expectations, i.e. by the following recursive formula:
$\mathcal{O}_{L}\left(\sigma_{V}\right):=\mathcal{O}\left(\sigma_{V}\right),\ \ \ \mathcal{O}_{\ell-1}\left(\sigma_{Q_{\ell-1}}\right):=\langle\mathcal{O}_{\ell}\left(\sigma_{Q_{\ell}}\right)\rangle_{\xi_{\ell}},\ \ \ \mathcal{O}_{0}=\langle\mathcal{O}\left(\sigma_{V}\right)\rangle_{\mu}.$ (54)
We iterate the recursion from the initial condition $\mathcal{O}_{L}$ until the last step $\mathcal{O}_{0}$ that correspond the thermal average of $\mathcal{O}$ according to $\mu$. Introducing the free energy functional
$f\left(\sigma_{V}\right):=-\frac{1}{N}\log Z\left(\sigma_{V}\right)$ (55)
that is a flat functional of the order parameter $p$, we can apply the formula before
$f\left(\sigma_{V}\right)=\sum_{\ell\leq L}p_{\ell}f_{\ell}\left(\sigma_{Q_{\ell-1}}\right),\ \ \ f_{\ell}\left(\sigma_{Q_{\ell-1}}\right):=-\frac{1}{|V_{\ell}|}\log Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)$ (56)
and write the actual free energy in terms of the flat functional
$Nf=-\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\,\rangle_{\mu}.$ (57)
The usual free energy principle is recovered as before by bringing the average respect to $\mu$ inside the exponential, introducing the functional
$\mathcal{F}^{*}\left(p\right):=N\langle\,f\left(\sigma_{V}\right)\rangle_{\mu}$ (58)
we obtain the upper bound to the free energy:
$Nf\leq\inf_{p\,\in\mathscr{P}(\Lambda)}\mathcal{F}^{*}\left(p\right)$ (59)
Similarly, for a quenched free energy functional
$N\langle f\rangle_{\eta}=-\langle\,\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\rangle_{\mu}\rangle_{\eta},$ (60)
we can bring the average respect to $\eta$ inside the logarithm
$\mathscr{G}^{*}\left(p\right):=-\log\,\langle\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\rangle_{\mu}\rangle_{\eta}$ (61)
and find a lower bound to the quenched free energy:
$\langle f\rangle_{\eta}\geq\sup_{p\,\in\mathscr{P}(\Lambda)}\mathscr{G}^{*}\left(p\right)$ (62)
We expect that this second variational principle is of the same kind of that appearing in the celebrated Parisi formula for the SK model [4] where, in fact, the prescription is to maximise the the energy functional respect to the order parameter.
2.2 Blanket internal structure
It has been shown [1] that the blankets possess an internal structure. As one can easily appreciate from Figure 1 of [1], is possible to identify at lest two main types of energy contributions, those corresponding to edges with both ends in $V_{\ell}$, that we call the core, and those with one edge in $V_{\ell}$ and one in $Q_{\ell-1}$, that we call interface:
$W_{\ell}=\hat{W}_{\ell}\cup\bar{W}_{\ell},\ \ \ \hat{W}_{\ell}:=V_{\ell}^{2},\ \ \ \bar{W}_{\ell}:=\left(V_{\ell}\,Q_{\ell-1}\right)\cup\left(Q_{\ell-1}V_{\ell}\right)$ (63)
Hereafter, we indicate with a hat on top if the quantity is relative to the core and with a bar if to the interface. This partition is also shown in Figures 4.1 and 4.2 of [3]. The effect on the Hamiltonian of this further partition is best understood by introducing the normalisation coefficients
$\hat{\gamma}_{\ell}:=g\left(V_{\ell}\right)/g\left(V\right),\ \ \ \bar{\gamma}_{\ell}:=g\left(Q_{\ell}\right)/g\left(V\right)$ (64)
Using these coefficients, that for densely connected models can be non–trivial, due to the dependence of $g\left(V\right)$ on the number of spins, the core of the blanket (or layer) is
$\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right):=\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)$ (65)
and is simply a smaller model multiplied by a normalisation. The interface is instead described by a very simple bipartite Hamiltonian, that is the scalar product between the spins of the blanket and the local fields:
$\bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right):=\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)$ (66)
The full blanket is then recovered by summing back these contributions
$H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right)+\bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)$ (67)
Notice that if we interpret $\ell$ as some sort of discrete time, then the the blanket structure is equivalent to the celebrated Transformer circuit [25], with the cores corresponding to the the attention modules [26]. Let conclude by studying the effects of this decomposition on the partition function: to properly explain this it will be convenient to introduce the partition function of the interface,
$\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right):=\prod_{i\in V_{\ell}}2\cosh\left[\bar{\gamma}_{\ell}h_{i}\left(\sigma_{Q_{\ell-1}}\right)\right]$ (68)
and the associated Gibbs measure (softmax)
$\bar{\xi}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right):=\frac{\exp\left[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)\right]}{\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)}$ (69)
Then, the average respect to the interface is given by the formula [2]
$\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\rangle_{\bar{\xi}_{\ell}}=\frac{1}{\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)}\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\left[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)\right]$ (70)
Concerning the core, the partition function is just a number as all spin degrees of freedom gets integrated out, the Gibbs average is then just a smaller model restricted to the spins of the blanket and rescaled by a factor $\hat{\gamma}_{\ell}$
$\hat{Z}_{\ell}:=\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\exp\,[-\hat{\gamma}_{\ell}\,H\left(\sigma_{V_{\ell}}\right)],\ \ \ \hat{\xi}_{\ell}\left(\sigma_{V_{\ell}}\right):=\frac{\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]}{\hat{Z}_{\ell}}$ (71)
the average is indicated like before
$\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\rangle_{\hat{\xi}_{\ell}}=\frac{1}{\hat{Z}_{\ell}}\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]$ (72)
Let compute the full blanket average. We start from the identity
$\exp\,[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)]=\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\,\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]$ (73)
and integrate the spin variables in two steps. Notice that the order in which the integrations are performed, although not influent on the final result, can give rise to different (equivalent) expressions. For example, if we integrate the core first
$\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\rangle_{\xi_{\ell}}=\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\hat{\xi}_{\ell}\left(\sigma_{V_{\ell}}\right)\frac{\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]}{\langle\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]\rangle_{\hat{\xi}_{\ell}}}=\\ =\frac{\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]\rangle_{\hat{\xi}_{\ell}}}{\langle\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]\rangle_{\hat{\xi}_{\ell}}}$ (74)
while integrating the interface first we find
$Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\exp\,[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)]=\\ =\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\,\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]=\\ =\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\bar{\xi}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]=\\ =\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)\langle\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\rangle_{\bar{\xi}_{\ell}}$ (75)
for the partition function and
$\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\rangle_{\xi_{\ell}}=\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\frac{\exp\,[-H_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)]}{Z_{\ell}\left(\sigma_{Q_{\ell-1}}\right)}=\\ =\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\frac{\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\,\exp\,[-\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)]}{\bar{Z}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)\langle\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\rangle_{\bar{\xi}_{\ell}}}=\\ =\sum_{\sigma_{V_{\ell}}\in\Omega^{V_{\ell}}}\bar{\xi}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)\frac{\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]}{\langle\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\rangle_{\bar{\xi}_{\ell}}}=\\ =\frac{\langle\mathcal{O}\left(\sigma_{V_{\ell}}\right)\,\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\rangle_{\bar{\xi}_{\ell}}}{\langle\exp\,[-\hat{\gamma}_{\ell}H\left(\sigma_{V_{\ell}}\right)]\rangle_{\bar{\xi}_{\ell}}}$ (76)
for the full blanket average. What is interesting in the first case in Eq. (74) (core first) is that the averaged quantities all depend on the core distributions. This is the approach that was actually taken in [1] to prove the dominance of the interfaces in the SK model and may be convenient in some special cases, but we now believe that the most useful expression is the second case in Eq. (76) (interface first), because the averages are done respect to the “interface” model [2] that is a one–body Hamiltonian, is universal and is generally much easier to control, as is equivalent to adding an external field.
3 Fully connected models
In [1, 2, 3] we showed that in fully connected models with reasonable coupling strength distribution the core can be ignored if the partition is fine enough.
In general, for finite dimensional models, or any other networks with irreducible core, we have that the ratio between the core and the interface converges to a constant
$\hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}\rightarrow\mathrm{const.}$ (77)
and both the cores and the interfaces contribute to the total energy of the blankets in a comparable way also for an infinitely refined partitions, ie. for $L\rightarrow\infty$. In this case the theory is nontrivial and non–mean field. This class includes eg. the Ising model in two and three dimension, polymers [22, 23, 24], random models like the Edaward-Anderson model, but also the Transformer architecture by Google [25, 26] the neural Lattice Field Theory (LFT) in Section 4 of [10], etc.
Then, we have two types of mean field models: the first is what we call the core-only (CO) model, is such that the interface can be ignored
$\hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}\rightarrow\infty,\ \ \ f\rightarrow\hat{f}$ (78)
and is the most typical among the mean–field approximations [27]. Practically, it consists in finding a partition (if any) in which the interfaces can be neglected. In such way the measure is just the product measure of the cores, the blankets are approximated as independent.
Finally, the second kind of mean–field approximation considered in [1, 2, 3] is the interface only (IO) model, where
$\hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}\rightarrow 0,\ \ \ f\rightarrow\bar{f}$ (79)
i.e., where the core contributions are overwhelmed by the interfaces, or where we can identify the blanket sequence in such way that the core is zero, e.g. like a purely feed–forward network, or the levels of the Caylay tree, if we want to consider a model with finite connectivity. As we shall see, this class includes the Curie–Weiss model and the SK model.
3.1 General theory of the IO model
Since the interface seems to have a quite universal scaling limit let construct a general theory of the IO model [3]. We will start from the IO Hamiltonian
$\bar{H}\left(\sigma_{V}\right):=\sum_{\ell\in\Lambda}\sigma_{V_{\ell}}\cdot\bar{\gamma}_{\ell}h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)$ (80)
If we arrange in decreasing order the field components we can further split $V_{\ell}$ into $L^{\prime}$ sub-blankets with approximately constant value. We call this partition $\mathscr{V}_{\ell}$
$V_{\ell}=\{V_{\ell\ell^{\prime}}\in\mathscr{V}_{\ell}:\,\ell^{\prime}\in\Lambda^{\prime}\},\ \ \ \Lambda^{\prime}:=\{1\leq\ell^{\prime}\leq L^{\prime}\}$ (81)
For simplicity hereafter we take $L^{\prime}=L$. We can identify two bounds, both dependent on the previous blankets but not on sub-blankets of the same blanket, and such that
$h_{\ell\ell^{\prime}}^{-}\left(\sigma_{Q_{\ell-1}}\right)\leq h_{i}\left(\sigma_{Q_{\ell-1}}\right)\leq h_{\ell\ell^{\prime}}^{+}\left(\sigma_{Q_{\ell-1}}\right),\ \ \ \forall i\in V_{\ell\ell^{\prime}}$ (82)
Most important, the gap shrinks proportionally to the inverse of $L^{2}$ [3]
$h_{\ell\ell^{\prime}}^{+}\left(\sigma_{Q_{\ell-1}}\right)-h_{\ell\ell^{\prime}}^{-}\left(\sigma_{Q_{\ell-1}}\right)=O\left(1/L^{2}\right)$ (83)
and is therefore converging to zero in the limit of infinite layers (blankets). Introducing some useful additional notation
$\hat{\gamma}_{\ell\ell^{\prime}}:=g\left(V_{\ell\ell^{\prime}}\right)/g\left(V\right),\ \ \ \hat{h}_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right):=\hat{\gamma}_{\ell\ell^{\prime}}h_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right)$ (84)
we immediately see that the Hamiltonian of the IO model satisfies
$\sum_{\ell\in\Lambda}\,\sum_{\ell^{\prime}\in\Lambda^{\prime}}\hat{h}_{\ell\ell^{\prime}}^{-}(\sigma_{Q_{\ell-1}})\,m(\sigma_{V_{\ell\ell^{\prime}}})\leq\frac{\bar{H}\left(\sigma_{V}\right)}{N}\leq\sum_{\ell\in\Lambda}\,\sum_{\ell^{\prime}\in\Lambda^{\prime}}\hat{h}_{\ell\ell^{\prime}}^{+}(\sigma_{Q_{\ell-1}})\,m(\sigma_{V_{\ell\ell^{\prime}}})$ (85)
and can be therefore approximated by energy functions that are supported by the magnetisation eigenstates, similarly to what conjectured by Mezard and Virasoro in [7] for the SK model. This is important, as it allows to work also in the thermodynamic limit using well known methods [28]. Most important, we found that under Gibbs measure the fields are bounded from above/below by the set of parameters [1, 2, 3]
$h^{\pm}\in\mathbb{R}^{\,L^{2}}$ (86)
in the thermodynamic limit, and that the sub-blanket average
$\lim_{N\rightarrow\infty}\langle\mathcal{O}\left(\sigma_{V_{\ell\ell^{\prime}}}\right)\rangle_{\bar{\xi}_{\ell\ell^{\prime}}}=\lim_{N\rightarrow\infty}\langle\mathcal{O}(\sigma_{V_{\ell\ell^{\prime}}})\rangle_{m_{\ell\ell^{\prime}}}$ (87)
will converge to an average over magnetisation eigenstates with given eigenvalue. This last step allows to actually compute the interface averages numerically as function of the parameters by applying the recursion in Eq. (54) for each layer and sub–layer.
Notice that the magnetisation eigenstates commute, i.e., the overlap matrix and the correlation matrix both converge to the square of the magnetisation in the thermodynamic limit [2]. The implications of this are explained in Section 2 of [3].
Also, in [2, 3] we showed that, at least in the small temperature limit, the IO model converges to a Generalised Random Energy Model (GREM) of the Derrida’s type [29, 30] due to a general phenomenon known as REM universality [31], and that this result can be extended to any temperature at least in case of Gaussian couplings [2].
Finally, notice that due to the new level of partitioning the order parameter evolved from a vector into a matrix, but then, notice that since there is no space structure only the levels matter, and the order parameter is ultimately the distribution of the entries of the parameter matrix.
3.2 Free energy functional
Now that we have a method to compute the averages it only remains to find a manageable expression for the free energy functional. We start from
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\left[\bar{\gamma}_{\ell}\,h_{i}\left(\sigma_{Q_{\ell-1}}\right)\right]$ (88)
This is the normalised expression for the free energy density of the blanket. Notice that for uniform models (i.e. when the order of the spins don’t matter) it is possible to make the parts of any size we want: for example, taking the finest
[…]
the coefficient are like before
$\hat{\gamma}_{i}:=g\left(V_{i}\right)/g\left(V\right),\ \ \ \bar{\gamma}_{i}:=g\left(Q_{i}\right)/g\left(V\right)$ (90)
but see the most elementary energy contribution is now composed by just a single spin coupled to an element of the coupling matrix
$h_{i}(\sigma_{j}):=H_{ij}\sigma_{j}$ (91)
The cavity field is then rewritten as follows
$\bar{\gamma}_{i}\,h_{i}(\sigma_{Q_{i-1}})=\sum_{j<i}\hat{\gamma}_{j}h_{i}(\sigma_{j})$ (92)
and by substituting in the definition we find
$\bar{f}_{i}\left(\sigma_{Q_{i-1}}\right)=-\log 2\cosh\left[\bar{\gamma}_{i}\,h_{i}(\sigma_{Q_{i-1}})\right]=-\log 2\cosh\,\sum_{j<i}\hat{\gamma}_{j}\,h_{i}(\sigma_{j}).$ (93)
In the thermodynamic limit the index is continuous [2], let
$i/|V_{\ell}|\rightarrow x\in\left[0,1\right],\ \ \ j/|V_{\ell}|\rightarrow y\in\left[0,1\right]$ (94)
the coefficients and the cavity fields converge to functions
$\bar{\gamma}_{i}\rightarrow\gamma\left(x\right),\ \ \ \hat{\gamma}_{j}\rightarrow d\gamma\left(y\right),\ \ \ h_{i}(\sigma_{j})\rightarrow h\left(x,y\right)$ (95)
putting together, the continuum limit of the interface is given by the integral [2]
[…]
then if the field is nonrandom we can simply apply the chain rule, while for fields with random coefficients the expression converges to an Ito integral that can be analysed by Ito calculus, see for example [32]. An alternative way to find the free energy functional for the recursion of Eq. (54) is the “Cavity Method” (CM), [4]. The CM is shown in the kernel context in the Section 6 of [3] for the SK model, using a mathematical approach originally developed by Aizenmann, Sims and Starr [33] (see also [1] or [34]).
3.3 Example 1: Curie–Weiss model
As first practical application we analyse the Curie-Weiss model [35], a fully connected model where the interaction strength is constant. The parameters are
$H_{ij}=\beta J_{0}/N,\ \ \ g\left(V\right)=N$ (97)
Since the prescription for the existence of the thermodynamic limit is that the components of the cavity fields stay finite as the system grows, then the cavity field must be normalised by a quantity that is linear in the number of spin. Then we must have the following form of the cavity field
$h_{i}\left(\sigma_{V}\right)=\frac{J_{0}}{N}\sum_{j\in V}\sigma_{j}=J_{0}\,m\left(\sigma_{V}\right)$ (98)
and the full Hamiltonian is just proportional to the square of the magnetisation operator
$H_{\mathrm{CW}}\left(\sigma_{V}\right):=\frac{J_{0}}{N}\sum_{i\in V}\sum_{j\in V}\sigma_{i}\sigma_{j}=NJ_{0}\,m\left(\sigma_{V}\right)^{2}$ (99)
Most important, we can explicitly compute the coefficients
$\hat{\gamma}_{\ell}=p_{\ell},\ \ \ \bar{\gamma}_{\ell}=q_{\ell-1},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=p_{\ell}/q_{\ell}\rightarrow 0$ (100)
and find, as predicted, that in the TL and infinite levels of partition the core energy
$\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right)=p_{\ell\,}H_{\mathrm{CW}}(\sigma_{V_{\ell}})=NJ_{0}\,p_{\ell}^{2}\,m\left(\sigma_{V_{\ell}}\right)^{2}$ (101)
becomes negligible in magnitude respect to the interface energy
$\bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=NJ_{0}\,q_{\ell-1}m\left(\sigma_{Q_{\ell-1}}\right)=NJ_{0}\,\sum_{\ell^{*}<\ell}p_{\ell^{*}}m\,(\sigma_{V_{\ell^{*}}})$ (102)
Notice in the last step we expanded the magnetisation states into the local magnetisations: we used a different superscript for the auxiliary variable $\ell^{*}$ to avoid confusion with the sub-index $\ell^{{}^{\prime}}$ before that has quite different meaning. The free energy functional of the blanket is then
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{\beta}\log 2\cosh\,[\,\beta J_{0}\,q_{\ell-1}m\,(\sigma_{Q_{\ell-1}})]=\\ =-\frac{1}{\beta}\log 2\cosh\,\beta J_{0}\sum_{\ell^{*}<\ell}p_{\ell^{\prime}}m\,(\sigma_{V_{\ell^{*}}})$ (103)
and we can explicitly compute the free energy density by applying the cascade of conditional expectations described in Eq. (54).
3.4 Example 2: Sherrington–Kirckpatrick model
Let now move to the glassy version of the CW model, that is the SK model. The model parameters are random this time
$H_{ij}=\beta J_{ij}/\sqrt{N},\ \ \ J_{ij}\sim\mathcal{N}\left(0,1\right),\ \ \ g\left(V\right)=1/\sqrt{N}$ (104)
the Hamiltonian of the (asymmetric) SK model is
$H_{\mathrm{SK}}\left(\sigma_{V}\right):=\frac{1}{\sqrt{N}}\sum_{i\in V}\sum_{j\in V}J_{ij}\sigma_{i}\sigma_{j}$ (105)
the normalisation is the square root of the number of spins
$h_{i}\left(\sigma_{V}\right)=\frac{1}{\sqrt{N}}\sum_{j\in V}J_{ij}\sigma_{j}$ (106)
this immediately imply the following from of the coefficients
$\hat{\gamma}_{\ell}=\sqrt{p_{\ell}},\ \ \ \bar{\gamma}_{\ell}=\sqrt{q_{\ell-1}},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=\sqrt{p_{\ell}/q_{\ell}}\rightarrow 0$ (107)
and by confronting the expressions of the core vs interface energies
$\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right)=\sqrt{p_{\ell}}\ H_{\mathrm{SK}}\left(\sigma_{V_{\ell}}\right),\ \ \ \bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=\sigma_{V_{\ell}}\cdot\sqrt{q_{\ell-1}}\ h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)$ (108)
we can see that the interface contribution is still dominant in the thermodynamic limit. Then, expanding the cavity field like before
$\sqrt{q_{\ell-1}}\ h_{i}(\sigma_{Q_{\ell-1}})=\sum_{\ell^{*}<\ell}\sqrt{p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (109)
we arrive to the free energy functional of the SK model
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sqrt{q_{\ell-1}}\,\,h_{i}(\sigma_{Q_{\ell-1}})=\\ =-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sum_{\ell^{*}<\ell}\sqrt{p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (110)
As said before, an alternative way to find the functional is the Cavity Method, [4, 33]: we showed a detailed derivation for the SK model in [1, 2, 3] (see also ref. [34] for the shortest derivation). Also, look at the Section 4, paragraph 4.2.1 of [10] for how to obtain the RSB theory as special (stationary) case of the kernel picture, within the larger context of Lattice Field Theories [3, 10, 11, 19, 20].
Acknowledgements
We are grateful to Pan Liming (USTC), Giampiero Bardella (Sapienza Università di Roma) and Nicola Kistler (Goethe University Frankfurt) for interesting discussions and suggestions. This research was partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 694925).